rite down the algorithm then use the numbers. greatest common divisor of the following pairs of a. 315, 825 b. 2091, 4807
Answer #1

O given two valuus be abl =@sg25 a-b) = (3 i 5,325 a= 315 hlowi Ly uging Eucl id ean Alaprim 825-315(2) t 195 54sC)30 45 = 35 40) 4- so, the gcd C2091 , 4804) = 1. I° int gcd Lint a, in-b) make a Swap Caib :︵ a mod b a - b end